OpenStudy (anonymous):

Instructions: Help please. Given that log 4= 0.6021, log 9= 0.9542 and log 12= 1.0792, evaluate each logarithm. Problem---> 20. log 0.00009

OpenStudy (tkhunny):

\(0.00009 = 9*10^{-5}\)

OpenStudy (anonymous):

why is that ?

OpenStudy (tkhunny):

Definition \(9 = 9*10^{0}\) \(0.9 = 9*10^{-1}\) \(0.09 = 9*10^{-2}\) \(0.009 = 9*10^{-3}\) \(0.0009 = 9*10^{-4}\) It's just the way it works.

OpenStudy (anonymous):

wouldn't it be log 10^-5 + log 9 ?

OpenStudy (anonymous):

then -5 + 0.9542 <-- (log 9) = 5.9 ?

OpenStudy (tkhunny):

??? What happened to the negative sign on the 5?

OpenStudy (anonymous):

its still there... my bad i calculated wrong , answer ? --> -4.0458 ?

OpenStudy (tkhunny):

That's better.

OpenStudy (anonymous):

that would be the answer ?

OpenStudy (tkhunny):

You were expecting something else? It looks good to me.

OpenStudy (anonymous):

no i just wanted to be reassured that, that was the correct answer.

OpenStudy (tkhunny):

There's a lot of that going around. Keep doing things well and you will develop more confidence in your own work.